Doreen Schmidt is a chemist. She needs to prepare 32 ounces of a 9% hydrochloric acid solution. Find the amount of 16% solution and the amount of 8% solution she should mix to get this solution
Answer by ankor@dixie-net.com(22740)

She needs to prepare 32 ounces of a 9% hydrochloric acid solution.
Find the amount of 16% solution and the amount of 8% solution she should mix to get this solution
:
let x = amt of 16% solution
the total is to be 32 oz, therefore;
(32-x) = amt of 8% solution
:
.16x + .08(32-x) = .09(32)
.16x + 2.56 - .08x = 2.88
.16x - .08x = 2.88 - 2.56
.08x = .32
x = .32/.08
x = 4 oz of 16% solution required
and
32 - 8 = 28 oz of 8% solution
:
:
:
Check this
.16(4) + .08(28) = .09(32)
.64 + 2.24 = 2.88